OpenStudy (anonymous):

How do i solve this trig equation? tan(3Θ)+1=0 where 0<=Θ<2π

OpenStudy (anonymous):

can you divide the 3 out of the tanΘ?

OpenStudy (mathmale):

There are several ways, but I think one of the easier to follow is to replace 3 theta temporarily with x and solve the resulting equation: tan x + 1

OpenStudy (anonymous):

would you have to change the restriction to \[0\le3\theta <6\Pi\]

OpenStudy (mathmale):

that is, tan x + 1 + 0. for what angle or angles, between 0 and 2pi, is tan x = -1?

OpenStudy (anonymous):

the reference angle for 1 is pi/4, and tan is negative is quadrants 2 and 4 so it is 3pi/4 and 7pi/8?

OpenStudy (mathmale):

would you have to change the restriction to 0≤3θ<6Π ? Let's solve the problem; after which I'd bet you could answer your own question.

OpenStudy (mathmale):

tan 3Pi/4 is indeed equal to -1. I'm not convinced that 7Pi/8 is a solution; have you checked it in tan x + 1 = 0?

OpenStudy (anonymous):

ohh i meant 7pi/4

OpenStudy (anonymous):

so then you set that equal to tan(3Θ) and divide by 3

OpenStudy (mathmale):

You probably remember that the period of the tangent function is pi (not 2pi as with the sine and cosine). Almost. Set 3pi/4 equal to 3 theta and solve for theta. Then do the same for 7pi/8. checking your own answer (value(s) of theta) will tell you whether you're right that 7pi/4 is a solution of tan x + 1 = 0.

OpenStudy (anonymous):

ok, thanks!

OpenStudy (mathmale):

My pleasure. Shall i assume we're done?

OpenStudy (anonymous):

yes!

OpenStudy (mathmale):

Good for you. Hope to work with you again.